OpenStudy (anonymous):

In 360 days the radioactivity of a sample decreases by 74 percent. what is the half life? In 360 days the radioactivity of a sample decreases by 74 percent. what is the half life? @Mathematics

OpenStudy (anonymous):

solve \[.74^t=\frac{1}{2}\] for t in one step

OpenStudy (anonymous):

im in calculus 2 not physics

OpenStudy (anonymous):

OH using y=Ce^kt

OpenStudy (anonymous):

you get \[t=\frac{\ln(\frac{1}{2})}{\ln(.74)}\]

OpenStudy (anonymous):

of course your answer will be in years, not days

OpenStudy (anonymous):

840?

OpenStudy (anonymous):

hold on i have to read carefully

OpenStudy (anonymous):

it says "decreases by 74%" not "has 74% of its initial value" so i was incorrect.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if it decreases by 74% it has only 26% of its initial value. so you have to solve \[.26^t=.5\] instead

OpenStudy (anonymous):

same idea though, get \[t=\frac{\ln(.5)}{\ln(.26)}\] and again your units are years. but if you want days that is easy enough

OpenStudy (anonymous):

i get 187.8 days rounded.

OpenStudy (anonymous):

it says that is wrong, but it wont give me the correct answer

OpenStudy (anonymous):

how many decimal places does it want?

OpenStudy (anonymous):

its not specific and i already tried all the decimals

OpenStudy (anonymous):

I GOT IT

OpenStudy (anonymous):

ok maybe try this one but i don't think it is right http://www.wolframalpha.com/input/?i=log%28.5%29%2Flog%28.74%29*365

OpenStudy (anonymous):

oh what was the problem?

OpenStudy (anonymous):

u multiply by the 360 not 365

OpenStudy (anonymous):

lordamercy. nothing like the ability to read. i guess i just assumed it was 365 when i read it.

OpenStudy (anonymous):

can you show me how you derived .26^t=.5 from the y=Ce^kt formula

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first off the C is unimportant, because you want the half life. so it doesn't matter what you start with, at the end you will have half of the original amount. in other words if i start with \[C\times .26^t=\frac{1}{2}C\] the first step will be to rewrite as \[.26^t=\frac{1}{2}\]

OpenStudy (anonymous):

as for the .26 i reason that if you lose 74% you retain 26% (not much to that, 100 - 74 = 26)

OpenStudy (anonymous):

ok so now the next part of the question asks How long will it take for a sample of 100mg to decay to 81 mg

OpenStudy (anonymous):

and therefore i don't need the \[Ce^kt\] formula which requires me to solve for k using logs etc. i just say that every year it is 26% of the previous year, so i use \[.26^t\] instead of \[e^{kt}\]

OpenStudy (anonymous):

ok now we have to be careful

OpenStudy (anonymous):

since i messed up with the year business. we need to use \[100\times .26^{\frac{t}{360}}=81\]

OpenStudy (anonymous):

ya that makes sense not because C is the initial

OpenStudy (anonymous):

divide by 100 get \[.26^{\frac{t}{360}}=.81\]

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

right, but i have to use an exponent of \[\frac{t}{360}\] because it decrease to 26% of its current amount every 26 days

OpenStudy (anonymous):

360 days rather

OpenStudy (anonymous):

solve as before \[\frac{t}{360}=\frac{\ln(.81)}{\ln(.26)}\] \[t=360\times \frac{\ln(.81)}{\ln(.26)}\]

OpenStudy (anonymous):

ya i got that, let me try to put in the answer

OpenStudy (anonymous):

yup got it, thanx alot

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i got a couple more question concerning decay and growth, so if you can just keep watch for my posts, i really appreciate your help

OpenStudy (anonymous):

no problem, but post quickly because i have to go soon enough

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

btw you can use \[Ce^{kt}\] if you like, but it requires finding k, and then when you use the decimal there will be rounding error

OpenStudy (anonymous):

you have 700 dollars in your bank account. Suppose your money is compounded every month at a rate of 0.5 percent per month. (a) How much do you have after t years.

OpenStudy (anonymous):

C=700 but how do i set up the rest of the equation